OpenStudy (anonymous):

the average of 8 people is 22 when the oldest people left the average decrease to 20 how old was the person who left the room???

OpenStudy (anonymous):

Total age of 8 people=8*22=176 Total age of 7 people=20*7=140 Eighth person is 36 years old